OpenStudy (anonymous):

Factor y2 + 5y + 6. (y + 1)(y + 6) (y + 2)(y + 3) (y + 2)(y + 4)

OpenStudy (anonymous):

y^2+5y+6=0 y^2+3y+2y+6=0 y(y+2)+2(y+3)=0 (y+2)(y+3)=0